OpenStudy (anonymous):

At a basketball game, a vender sold a combined total of 240 sodas and hot dogs. The number of hot dogs sold was 52less than the number of sodas sold. Find the number of sodas sold and the number of hot dogs sold

OpenStudy (anonymous):

Total Number sold = x + (x-52) where x is the number of sodas sold 240 = x + x - 52 292 = 2x x = 146 146 Sodas sold, 94

OpenStudy (anonymous):

94 hot dogs sold*

OpenStudy (anonymous):

cool thanks